40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 22 19:53:01 2013, Jenne, Update, Locking, PR-flat cavity status - not locked 
    Reply  Wed Jan 23 19:05:16 2013, Jenne, Update, Locking, PR-flat cavity status - locks! 
       Reply  Wed Jan 23 20:46:46 2013, Zen Master, Update, Locking, PR-flat cavity status - locks! yinYang.png
Message ID: 7931     Entry time: Wed Jan 23 19:05:16 2013     In reply to: 7928     Reply to this: 7934
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PR-flat cavity status - locks! 

Status update

I (with help from Q) have redone the POP path on the ITMX table.  1" iris is a little too small, so I took it out.  2" lens moved to be centered on POP beam.  2" Y1 didn't need moving.  Straight refl from the 2" Y1 was aligned on to a PDA10CS (set to 70dB). This PD is blocking the usual POP55 diode.  BS which sends beam to camera was moved to allow room for the new temp DC PD.  Refl from this BS goes to the POP camera, which was moved so that the POP beam takes up most of the camera.  BS that would normally take half of the camera's beam and send it to POP22 (Thorlabs PD) is removed, so no beam to POP22.

Also, I have taken the output of the PDA10CS and hijacked the "POP110" heliax cable.  This was connected to this Thorlabs PD which is used as POP22.  (Kiwamu and I had long-term borrowed the 110 demod board for an AS 110 diode, so the "POP110" heliax was really only serving POP22.) There are yellow labels on the new temp and old regular cables, so we can undo my hack.  Similarly, on the other end of the heliax at the LSC rack, I have taken the heliax's output and sent it to the POPDC input on the whitening board.  Thus, the regular POPDC SMA cable is unplugged, but labeled again with big yellow labels.

In other news - the PR-flat cavity locks!!! 

Koji and I coarsely rotated the REFL11 phase such that the signal is predominantly in the I phase.  We set the LSC input matrix to use REFL11I for PRCL, and the output matrix is set to actuate on PRM.  Then we set the gain to -0.005, and it locked!!!!

 

EDIT:  I turned back on the PRM oplev (after Manasa aligned it and redid the out-of-vac oplev layout a bit), and the motion of the cavity is slightly reduced, although there's still a lot going on.  The cavity is vaguely well aligned, although it's time to go make sure that the beams are still on the REFL and TRANS PDs.  However, it's dinner time.

ELOG V3.1.3-